(x^2+1)^(1/2) + 2(x^2+1)^(-1/2)
factor out (x^2+1)^(-1/2).
Please show the steps to factor completly

Answer by stanbon(75887)

(x^2+1)^(1/2) + 2(x^2+1)^(-1/2)
factor out (x^2+1)^(-1/2).
Please show the steps to factor completly
----------------------
Divide each term by that common factor to get:
---
(x^2+1)^[(1/2)-(-1/2)] + 2(x^2+1)^[(-1/2)-(-1/2)]
-----
= (x^2+1)^1 + 2(x^2+1)^0
---
= x^2+1 + 2*1
---
= x^2+3
============
Cheers,
Stan H.